OpenStudy (anonymous):

Solve for x: x^2-1=x+1 A. -1 B. -2 C. 2 & -1 D. -1 & -2

OpenStudy (anonymous):

C.2&-1

OpenStudy (anonymous):

some thing is wrong with this question x^2-1=x+1 is the same as x^1=x+1 now any number to the power of 1 is its self. so how must it be equal seeing that 1 is added to x

OpenStudy (anonymous):

It would prob. make more sence to you if i put (x^2)-1=x+1. would it not be the same answer this way?

OpenStudy (anonymous):

rnoy is right. substitute your choices into the question and see which one solves it.

OpenStudy (anonymous):

Ok thank you!